OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

hero (hero):

1. First find the equation of the line by inserting each point separately into the formula y = mx + b to get: 3 = -5m + b -3 = m + b 2. Isolate b in each equation to get: b = 3 + 5m b = -3 - m 3. Set b = b, then solve for m: 3 + 5m = -3 - m 5m + m = -3 - 3 6m = -6 m = -6/6 m = -1 since m represents the slope, the slope is -1. Now, substitute m = -1 back into one of the original formulas to find b: b = 3 + 5(-1) b = 3 - 5 b = 8 So now we know the equation of the original line: y = -x + 8

hero (hero):

Now, to find the x-intercept, just set y = 0, and solve for x.

OpenStudy (anonymous):

ok

hero (hero):

Let me know what you get for x.

OpenStudy (anonymous):

wait what do mean by set y = 0, and solve for x ?

hero (hero):

yes

hero (hero):

y = -x + 8

hero (hero):

Set y = 0, and solve for x

OpenStudy (anonymous):

is it -1

OpenStudy (anonymous):

is that right ?

hero (hero):

If you set y to 0 you will have 0 = -x + 8 Then after you add x to both sides, you will get x = 8

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks

hero (hero):

I'm trying to figure out why you couldn't do that on your own.

hero (hero):

Do you even know how to solve algebraic equations?

OpenStudy (anonymous):

yes

hero (hero):

Well, so far, unfortunately, you haven't convinced me.

OpenStudy (anonymous):

i head went blank for a second, i just forgot.